Grade 12th pass9 grade science

Two cars started moving with initial velocities u and 2u. For the same deceleration, their respective stopping distances are in the ratio

Harshit Singh
4 Years agoGrade 12th pass

1 Answer

Yash Patari
4 Years ago
DEAR STUDENT

Explanation:

using V^2 = U^2 - 2*a*s

s = v^2 / (2*a)

their respective stopping distances are in the ratio

1:4

THANKS